OpenStudy (anonymous):

Given: P(8, 4) and Q(-1, 9). Write the component form of vector PQ. Show all work.

OpenStudy (anonymous):

PQ means the multiplication of P and Q

OpenStudy (anonymous):

Questioner has stipulated that it means the vector from P to Q (often shown with a directional arrow on top), so that's fine. Admittedly, questioner has not stipulated that P and Q are to be considered vectors rather than points, this can be inferred from the context.

OpenStudy (anonymous):

PQ = OQ - OP = (-8,5) = -8i + 5j

OpenStudy (anonymous):

OP is a vector from (0,0) to (8,4). or 8i+4j OQ is a vector from (0,0) to (-1,9). or -i+9j PQ= OQ-OP = (-i+9j) - (8i+4j) = (-9i+5j)

OpenStudy (amistre64):

<-9,5>

OpenStudy (anonymous):

thank you amistre:)

OpenStudy (amistre64):

:)

OpenStudy (amistre64):

"component" form is might be off; but thats how i read the question.

OpenStudy (anonymous):

Dear amistre,, I need your guidance about programming. could u please help me?

OpenStudy (anonymous):

yeah, component form includes the unit vectors

OpenStudy (anonymous):

I wanna know which parts of math are moer important in CS?

OpenStudy (anonymous):

discrete math usually (not always).

OpenStudy (anonymous):

and other parts like game theory or geometry or ...

OpenStudy (amistre64):

CS is computer science?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

the discrete stuff is good to know, and easy to program for; but all of the maths are important in CS, its mathmatical models that allow us to use and abuse the computers. I cant say for sure since I dont know what all CS entails tho

OpenStudy (anonymous):

thanks a lot:)

OpenStudy (amistre64):

youre welcome :) I find the more math I know, the easier I can make programs ....

OpenStudy (anonymous):

that's amazing. i said -1 - 8 was -8. foolish.

OpenStudy (amistre64):

:) typos happen typing and mathing are two seperate brain functions and tend to get crossed